OpenStudy (anonymous):

Solve the equation x − 3 = 8

OpenStudy (anonymous):

x − 3 = 8 x − 3+3 = 8+3 (adding 3 both sides) x=11

OpenStudy (anonymous):

To get x alone you must add 3 to both sides giving 8+3=11